# DIRECT: Deep Active Learning under Imbalance and Label Noise

### v2

@@ 4.1 An 1-D Reduction to Agnostic Active Learning
        Estimate separation threshold
$\widehat{j}^{k}\leftarrow$ $\operatorname*{arg\,max}_{j}(|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|$
|$-|\{y_{(i)}=1:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$
and break ties by choosing the index closest to $\frac{N}{2}$. Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
@@
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{q_{(i)}\}_{i=1}^{N}$.
  Define: For examples in the labeled set, we let $y_{(i)}$ denote the label of the example corresponding to $q_{(i)}$.
  Initialize: Version space: shortest segment of indices $[I,J]$, with $q_{(I)},q_{(J)}$ $\forall i\leq I,q_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,q_{(j)}\in L\implies y_{(j)}\neq k$.
@@
     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $q_{(I)}$, …, $q_{(J)}$ for annotation and insert to $L$.
     Find $[I^{\prime},J^{\prime}]\subset[I,J]$ such that
$I^{\prime},J^{\prime}=\operatorname*{arg\,min}_{i,j:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}\quad\text{where }$
^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
     Update version space $I,J\leftarrow I^{\prime},J^{\prime}$.
@@
4.2 Algorithm
We are now ready to state our algorithm as shown in Algorithm 1. The algorithm follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. Our set of labeled examples at anytime are maintained in $L_{t}$.
To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2) on the set of labeled examples. Our method is motivated by the fixed-budget algorithm for active classification, ACED (Katz-Samuels et al., 2021). Intuitively, in the non-noisy or separable setting, where the loss of $\mathcal{L}(h_{j^{\star}})=0$, we could run a binary search procedure to learn $j^{\star}$. However, we can’t assume that $j^{\star}$ perfectly separates the classes. ACED circumvents this by maintaining a version space of threshold classifiers, namely an interval $[I,J]$. In each of the $b/B_{\text{parallel}}$ rounds, $B_{\text{parallel}}$ samples are taken and appended to the set of labeled examples $L$. Then, an interval $I^{\prime},J^{\prime}$ is computed, whose length is a factor of $1/c$ smaller than the current $I,J$ of low-loss classifiers. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left. The fixed confidence (or alternatively $\delta$-PAC) analogue of ACED is known to match instance optimal lower bounds for active classification up to logarithmic factors. Under benign noise conditions, the probability of misidentifying the optimal separation threshold decays exponentially w.r.t. budget $b$.
To address batch labeling, we note the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel. In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. As discussed above, $B_{\text{parallel}}\ll B_{\text{train}}$ in practice. Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
@@ 5 Experiments
    Fine-tuning ResNet-18 with vanilla supervised training on imbalanced datasets similar to Zhang et al. (2022).
@@
    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2023a).
In Section 5.1, we document these setups in detail. For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.2, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.3, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix B for all results.
5.1 Experiment Setup
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022). Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014). The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library. To address data imbalance, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019).
LabelBench Experiments. Proposed by Zhang et al. (2023a), LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021). The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021). Similar to the ResNet-18 experiments, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
Name # Classes $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
Imb SVHN $2$ $73257$0724.0724.0724.0724
Imb SVHN $3$ $54448$2546.2546.2546.2546
PathMNIST $2$ $89996$1166.1166.1166.1166
FMoW $62$ $76863$0049.0049.0049.0049
iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
@@ 5.1 Experiment Setup
(f) Imbalanced SVHN, two classes, number of labeled examples in minority class
Figure 3: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting. (a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework. $B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2. $B_{\text{parallel}}=1$ is equivalent with sychornous labeling. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
@@ 5.2 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty. For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results. In particular, we compared against BADGE, Margin Sampling (Balcan et al., 2006; Kremer et al., 2014), CORESET and GALAXY.
@@
As highlighted in Figures 1(a) and 3(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 3(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 3(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 3(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
@@
We would also like to highlight the ability to handle batch labeling. Across our experiments, we see DIRECT outperforms with different amounts of parallel annotation ($B_{\text{parallel}}=1,5$ and $20$), indicating its general effectiveness. This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$. On Figures 1(a) and 4(a), we see that DIRECT outperforms GALAXY with synchronous labeling. Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
@@
Figure 4: Performance of DIRECT against baseline algorithms under $10\%$ label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
@@
Figure 5: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
@@ 5.3 Experiments under Imbalance and Label Noise
Different Levels of Label Noise As shown in Figures 5(a)-(c) and 1(b), we observe the results on imbalanced CIFAR-100 with two classes across numerous levels of label noise, with $0\%$, $10\%$, $15\%$ and $20\%$ respectively. In fact, the noiseless experiment in Figure 5(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy. However, we see DIRECT becomes more label-efficient under the more practical label noise settings. It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling. In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
@@ 6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel. Studying asynchronous labeling could a natural extension of our work. We think a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
@@ Appendix A Equivalent Objective
Appendix B All Results
B.1 Noiseless Results under Imbalance

### v3

@@ 4.1 An 1-D Reduction to Agnostic Active Learning
        Estimate separation threshold (break ties by choosing the index closest to $\frac{N}{2}$):$\quad\widehat{j}^{k}\leftarrow\operatorname*{arg\,max}_{j}(|\{y_{(i)}=k:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|-|\{y_{(i)}\neq k:x_{(i)}\in L_{t}\,\text{and}\,i\leq j\}|)$. Annotate $b$ unlabeled examples with sorted indices closest to $\widehat{j}^{k}$ and insert to $L_{t}$.
@@
4.2 Algorithm
We are now ready to state our algorithm DIRECT as shown in Algorithm 1. Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. We spend half each round’s budget for both phases.
During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run a agnostic active learning procedure for the corresponding 1-D class-k v.s. rest reduction. Our agnostic active learning subroutine is formally outlined in Algorithm 2, VReduce. Our method corresponds to the fixed-budget ACED Katz-Samuels et al. (2021) algorithm, adopted to the threshold classifier scenario. To explain the intuition about ACED in this scenario, we first note that in the separable case with $\mathcal{L}(h_{j^{\star}})=0$, we could simply run a bisection procedure to learn $j_{k}^{\star}$, the optimal separation threshold for the $k$-th class. However, as $j^{\star}$ cannot perfectly separate the classes, ACED circumvents this by maintaining a version space of possible thresholds, namely the interval $[I,J]$. Statistically with high likelihood, the optimal separation threshold $j_{k}^{\star}$ lies in this version space of $[I,J]$. During each of the $m$ rounds of VReduce, $B_{\text{parallel}}$ samples are annotated and the version space’s length is shrank by a factor of $1/c$. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left. The second phase of DIRECT simply annotates examples closest to each optimal separation threshold, aiming to annotate a class-balanced and uncertain examples.
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels with sorted uncertainty $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (unlabeled $y_{(i)}^{k}$ hidden to learner).
  Initialize: Version space as the shortest segment of indices $[I,J]$ such that: for each labeled example $x_{(i)}$ with $i\leq I$, $y_{(i)}=k$, and for each labeled example $x_{(j)}$ with $j\geq J$, $y_{(j)}\neq k$.
@@ 4.2 Algorithm
     Uniformly sample $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}^{k}$, …, $x_{(J)}^{k}$ for annotation and insert to $L$.
     For each $0\leq s\leq N$, compute ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
     Shrink version space by $\frac{1}{c}$: $I,J\leftarrow\operatorname*{arg\,min}_{i,j\in[I,J]:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
@@
To address batch labeling, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. We note that, in practice, the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel, i.e., $B_{\text{parallel}}\ll B_{\text{train}}$. Lastly, as will be discussed in Section 6, our algorithm can also be modified for asynchronous labeling.
Theoretical Comparison with GALAXY. As mentioned in Section 3.2, GALAXY’s graph-based approach aims to identify all cuts and sample examples around all of them equally. On the other hand, DIRECT aims to identify only the separation threshold and sample around it, which is superior as we have argued before and shown in our results. We now present a more theoretical comparison. As we show in Appendix A, the graph-based approach in GALAXY will identify and annotate around at least one more cut in addition to the optimal separation threshold, with probability at least $1-\exp(-b\log(\frac{1}{1-\eta})/2)$. Here, $b$ is the budget of a single round of annotation and $\eta$ is the label noise ratio (see Appendix A for more details). This implies, when the budget $b$ is large, GALAXY will likely annotate around unnecessary cuts. This is in contrast with the agnostic active learning approach we take in DIRECT, where as shown by Katz-Samuels et al. (2021), the probability of misidentifying the optimal separation threshold decays exponentially w.r.t. budget $b$. In other words, with a large budget $b$, with high likelihood, DIRECT will focus its annotation around the optimal separation threshold. Lastly, time complexity analysis in Appendix C shows DIRECT’s superior speed compared to BADGE and GALAXY.
@@ 5 Experiments
    Supervised fine-tuning of ResNet-18 on imbalanced datasets similar to Zhang et al. (2022).
@@
    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2024a).
For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.1, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.2, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix E for complete results.
Experiment Setups. Our experiments utilize $10$ imbalanced datasets derived from popular computer vision datasets. For the ResNet experiments, we utilize imbalanced versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets. For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021) datasets. We refer the readers to Appendix D for more details on our experiment setups.
@@
(f) Imbalanced SVHN, two classes, # of minority labeled examples
Figure 4: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting. (a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework. $B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2. $B_{\text{parallel}}=1$ is equivalent with sychornous labeling. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
@@ 5.1 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty. For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results, including BADGE, Margin Sampling, CORESET and GALAXY.
@@
As highlighted in Figures 1(a) and 4(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 4(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 4(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 4(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
@@
We would also like to highlight the ability to handle batch labeling. Across our experiments, we see DIRECT outperforms with different amounts of parallel annotation ($B_{\text{parallel}}=1,5$ and $20$), indicating its general effectiveness. This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$. On Figures 1(a) and 5(a), we see that DIRECT outperforms GALAXY with synchronous labeling. Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
@@
Figure 5: Performance of DIRECT against baseline algorithms under $10\%$ label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
@@
Figure 6: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
@@ 5.2 Experiments under Imbalance and Label Noise
Different Levels of Label Noise As shown in Figures 6(a)-(c) and 1(b), we observe the results on imbalanced CIFAR-100 with two classes across numerous levels of label noise, with $0\%$, $10\%$, $15\%$ and $20\%$ respectively. In fact, the noiseless experiment in Figure 6(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy. However, we see DIRECT becomes more label-efficient under label noise. It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling. In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
@@ 6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel. Studying asynchronous labeling could be a natural extension of our work. A potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators. Lastly, we refer the readers to Appendix F for our impact statement.
@@ Appendix A Equivalent Objective
Appendix B Theoretical Analysis
In this section, we analyze the performance of GALAXY under random label noise and show the probability of identifying and sampling around additional cuts increases as more examples are labeled. This is in contrast to the DIRECT’s agnostic active learning approach, where the probability of identifying and sampling around only the optimal separation threshold decays exponentially in the number of labeling budget.
Specifically, under the binary classification scenario, one is given a sorted list of $N$ examples $\{x_{(i)}\}_{i=1}^{N}$, with ground truth labels $y_{(1)}^{\star}=y_{(2)}^{\star}=...=y_{(N_{1})}^{\star}=1$ and $y_{(N_{1}+1)}^{\star}=...=y_{(N_{1}+N_{2})}^{\star}=2$, where $N_{1}+N_{2}=N$. Under uniform i.i.d. label noise with noise ratio $\eta>0$, the observed labels are denoted as $\{y_{(i)}\}_{i=1}^{N}$, where $\mathbb{P}(y_{(i)}\neq y_{(i)}^{\star})=\eta$. In other words, the observed label is flipped with probability $\eta$.
Theorem B.1.
Given a budget of $b>2\log N$, let $M_{b}$ be the random variable denoting number of identified cuts in addition to the optimal separation threshold by one round of GALAXY. We must have $\mathbb{P}(M_{b}\geq 1)\geq 1-\exp(-b\log(\frac{1}{1-\eta})/2)$, implying GALAXY samples around at least one more cut in addition to the optimal separation threshold with high probability.
Proof.
In the perfect scenario where GALAXY does not receive any corrupted labels, it would use $\log N$ budget with bisection to find the optimal separation threshold and annotate around it. However, within the first $\frac{b}{2}$ annotations, whenever GALAXY receives a corrupted label, it will identify a cut in addition to the optimal separation threshold, i.e., $M_{b}\geq 1$. Therefore, the probability of $M_{b}\geq 1$ is greater than the probability of receiving at least one corrupted labels in the first $\frac{b}{2}$ annotations. With simple probability bound, we can show that
$\mathbb{P}(M_{b}\geq 1)>1-(1-\eta)^{b/2}=1-\exp(b\log(1-\eta)/2)=1-\exp(-b\log(\frac{1}{1-\eta})/2).$
As the theorem suggests, when $b$ is large, GALAXY will identify and annotate around at least one additional cut with high probability.
Appendix C Time Complexity
The computation complexity for each batch of DIRECT is $O(KN\log(N)+B_{\text{train}}N)$ for data selection plus the training and inference costs of the neural network. $O(KN\log(N))$ comes from sorting examples by their margin scores for each class and $O(B_{\text{train}}N)$ is the cost for running Algorithm 2 for $O(B_{\text{train}})$ iterations. Each iteration of Algorithm 2 only costs $O(N)$ time as we can efficiently solve the objective by cumulative sums. We note that the cost associated with neural network training and inference is always the dominating factor.
For comparisons, BADGE has time complexity $O(B_{\text{train}}N(K+D))$, significantly more expensive than DIRECT, with $D$ denotes the dimensionality of the penultimate layer features. In addition, GALAXY has computational complexity of $O(KN\log(N))+B_{\text{train}}KN$, also more expensive than DIRECT. In all of our experiments, both BADGE and GALAXY indeed is slower than DIRECT. We further note that the time complexity factor of $K$ in DIRECT can be easily parallelized by conducting the $K$ sorting procedures on different CPU cores.
In total, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
Appendix D Experiment Setup
Name $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
Imb SVHN $2$ $73257$0724.0724.0724.0724
Imb SVHN $3$ $54448$2546.2546.2546.2546
PathMNIST $2$ $89996$1166.1166.1166.1166
FMoW $62$ $76863$0049.0049.0049.0049
iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature [Ash et al., 2019, Zhang et al., 2022]. Our experiment setup utilizes the CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009], SVHN [Netzer et al., 2011] and PathMNIST [Yang et al., 2021] image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer [Kingma and Ba, 2014]. The ResNet-18 model [He et al., 2016] is pretrained on ImageNet [Deng et al., 2009] from the PyTorch library. To address data imbalance, for all algorithms, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training [Müller et al., 2019] for all algorithms.
LabelBench Experiments. Proposed by Zhang et al. [2024a], LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model [Radford et al., 2021]. The framework also utilizes semi-supervised learning method FlexMatch [Zhang et al., 2021] to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW [Christie et al., 2018] and iWildcam [Beery et al., 2021]. Similar to the ResNet-18 experiments, for all algorithms, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
Appendix E All Results
E.1 Noiseless Results under Imbalance